Question : in any triangle ∆ABC find the Least value of (sin^2A +SinA + 1)÷(SinA)

Grade:11

1 Answers

Manan Varma
7 years ago
The answer is 3. See [sin^2(A)+sin(A)+1]÷sin(A)=sin(A)+[1/sin(A)]+1We know that the minimum value sum of any number with it`s inverse is 2. Sosin(A)+[1/sin(A)] min value is 2So least value of the given expression is 2+1=3So the answer is 3.